Use the table to find the value that is three standard deviations above and below the mean. (Enter exact numbers as integers, fractions, or decimals.) 

Baseball PlayerBatting AverageTeam Batting AverageTeam Standard Deviation
Fredo0.1580.1160.012
Kart0.1770.1690.015

(a) above the mean 

Fredo's Team  _______ 

Kart's Team _______ 

(b) below the mean 

Fredo's Team _______ 

Karl's Team _______ 
Answer #1

Where is the mean and is standard deviation.

Since we are given teams's mean and standard deviations we simply add or subtract to find the below or above the mean.

Given

Fredo's team batting average = 0.116

Fredo's team standard deviation = 0.012

Karl's team batting average = 0.169

Karl's team standard deviation = 0.015

a.) Above the mean

Fredos team = 0.116 + 3*0.012 = 0.152

Karl's team = 0.169 + 3*0.015 = 0.214

b.) Below the mean

Fredo's team = 0.116 - 3*0.012 = 0.08

Karl's team = 0.169 - 3*0.015 = 0.124
Answer #2

ANSWER :


a. 3 SD above mean :


Fredo’s team  : 0.116 + 3*0.012 = 0.152 


Karl’s team : 0.169 + 3*0.015 = 0.214 


b. 3 SD below mean :


Fredo’s team  : 0.116 - 3*0.012 = 0.080


Karl’s team : 0.169 - 3*0.015 = 0.124 

answered by: Tulsiram Garg